# DG Singular equivalence and singular locus

### v5

@@ 4. Proof of the main result
Retain the settings in Proposition 3.11 and Lemma 3.12. Since the algebra morphism $\iota_{I}$ gives the $R/I$-module structure on $\mathrm{HH}_{sg}^{0}(R)$, the image of $\iota_{I}$ is contained in the central $Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}$ of $\mathrm{HH}_{sg}^{0}(R)$. Now, let
$(\iota_{I})_{\natural}:R/I\rightarrow Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}.$
be the induced map. By Lemma 3.12, it is clear that $(\iota_{I})_{\natural}$ is a finite morphism of Noetherian rings, i.e., $Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}$ is a finitely generated $R/I$-module. Hence, $Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}$ is also a Noetherian ring. By Proposition 3.11, we know that $(\iota_{I})_{\natural}$ is also an injection. It follows that there exists a surjective morphism of schemes:
$(\tilde{\iota}_{I})_{\natural}:\mathrm{Spec}\left(Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}\right)\twoheadrightarrow\mathrm{Spec}(R/I)$
given by $(\iota_{I})_{\natural}$. Moreover, $(\tilde{\iota}_{I})_{\natural}$ is also a finite morphism between schemes since $(\iota_{I})_{\natural}$ is a finite morphism of Noetherian rings. We thus completed the proof. ∎
@@
Again retain the settings in Proposition 3.11 and Lemma 3.12. Since $(\tilde{\iota}_{I})_{\natural}$ is a finite morphism between schemes, $(\tilde{\iota}_{I})_{\natural}$ has finite fibers, i.e., any one of its fibers has finite points. It implies that
$\mathrm{dim}\big{(}\mathrm{Spec}(R/I)\big{)}=\mathrm{dim}\left(\mathrm{Spec}\left(Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}\right)\right)$
since $(\tilde{\iota}_{I})_{\natural}$ is surjective. Thus the dimension of singular locous of $\mathrm{Spec}(R)$ is equal to the Krull dimension of $Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}$.
@@
$Z\big{(}\mathrm{HH}_{sg}^{0}(R_{1})\big{)}\cong Z\big{(}\mathrm{HH}_{sg}^{0}(R_{2})\big{)}$
@@ Acknowledgements.
We would like to thank Xiaojun Chen and Huijun Fan for their encouragement, support and suggestions. Moreover, we thank Martin Kalck for pointing out an error in the previous version of this paper.

### v6

@@ 4. Proof of the main result
Keep the settings in Proposition 3.11 and Lemma 3.12. Since the algebraic homomorphism $\iota_{I}$ gives the $R/I$-module structure on $\mathrm{HH}_{sg}^{0}(R)_{red}$, the image of $\iota_{I}$ is contained in the center $Z\big{(}\mathrm{HH}_{sg}^{0}(R)_{red}\big{)}$ of $\mathrm{HH}_{sg}^{0}(R)_{red}$. Note that $\mathrm{HH}_{sg}^{\bullet}(R)$ is a Gerstenhaber algebra, hence $\mathrm{HH}_{sg}^{0}(R)$ is a commutative subalgebra of $\mathrm{HH}_{sg}^{\bullet}(R)$ endowed with the cup product [26]. Therefore, $Z(\mathrm{HH}_{sg}^{0}(R)_{red})=\mathrm{HH}_{sg}^{0}(R)_{red}$. Now, we have a homomorphism between commutative rings
$\iota_{I}:R/I\hookrightarrow\mathrm{HH}_{sg}^{0}(R)_{red}.$
By Lemma 3.12, it is clear that $\iota_{I}$ is a finite morphism between commutative rings, i.e., $\mathrm{HH}_{sg}^{0}(R)_{red}$ is a finitely generated $R/I$-module. Hence, $\mathrm{HH}_{sg}^{0}(R)_{red}$ is also a Noetherian ring. Moreover, by Proposition 3.11, $\iota_{I}$ is also a finite injection. Combining above arguments, it follows that there exists a surjective morphism between schemes:
$(\iota_{I})_{\natural}:\mathrm{Spec}\left(\mathrm{HH}_{sg}^{0}(R)_{red}\right)\twoheadrightarrow\mathrm{Spec}(R/I)$
given by $\iota_{I}$. Furthermore, $(\iota_{I})_{\natural}$ is also a finite morphism between schemes by the finiteness property of $\iota_{I}$. Then we complete the proof. ∎
@@
Again keep the settings in Proposition 3.11 and Lemma 3.12. Since $(\iota_{I})_{\natural}$ is a finite morphism between schemes, $(\iota_{I})_{\natural}$ has finite fibers, i.e., any one of its fibers has finite points. It implies that
$\mathrm{dim}\big{(}\mathrm{Spec}(R/I)\big{)}=\mathrm{dim}\left(\mathrm{Spec}\left(\mathrm{HH}_{sg}^{0}(R)_{red}\right)\right)$
since $(\iota_{I})_{\natural}$ is surjective. Thus the dimension of singular locous of $\mathrm{Spec}(R)$ is equal to the Krull dimension of $\mathrm{HH}_{sg}^{0}(R)_{red}$.
@@
$\mathrm{HH}_{sg}^{0}(R_{1})_{red}\cong\mathrm{HH}_{sg}^{0}(R_{2})_{red}$
@@ Acknowledgements.
We would like to thank Xiaojun Chen and Huijun Fan for their encouragement, support and suggestions. Moreover, we thank Martin Kalck for pointing out an error in the previous version of this paper. This paper is supported by the National Natural Science Foundation of China (Grant No. 12401050).